2m^3 of soil containing 35% sand was mixed into 6m^3 of soil containing 15% sand. What is the sand content of the mixture?
Answer by Alan3354(69443)

2m^3 of soil containing 35% sand was mixed into 6m^3 of soil containing 15% sand. What is the sand content of the mixture?
--------------
2*0.35 = 0.7 cm of sand
6*0.15 = 0.9 cm of sand
--------------------
--> 1.6 m^3 of sand in 8 m^3
1.6/8 = 0.2 = 20% sand